The supplement of a angle is 4 more than 2 times it's complement. Find the measure of the angle.
Answer by ankor@dixie-net.com(12692)

The supplement of a angle is 4 more than 2 times it's complement. Find the measure of the angle.
:
Let x = the angle
:
(180-x) = it's supplement
and
(90-x) = it's complement
:
the equation of the given statement:
(180-x) = 2(90-x) + 4
:
180 - x = 180 - 2x + 4
:
-x + 2x = 180 - 180 + 4
:
x = 4
:
:
Check solution in original statement:
(180-4) = 2(90-4) + 4
176 = 2(86) + 4
176 = 172 + 4; confirms the solution of x=4